Scenario: I'm interested in whether there is a relationship between Team (A vs. B) and Outcome (Good vs. Bad). Below are the data. Test the null hypothesis that the categories are independent. α= .05.

Team A Team B
Good Outcome

fo = 35

fe=

fo = 55

fe=

Bad Outcome

fo = 15

fe=

fo = 45

fe=

The expected frequency, fe , for the Team A X Good Outcome category = _____

20

90

30

40

The Chi Square test statistic = _____

3.13

1.25

35.14

7.55

degrees of freedom, df = ______, and the Chi Square critical boundary = ______

1; 3.84

15; 25.00

3; 7.81

2; 5.99

What is the correct result based on the data?

Retain H0; Not Significant; categories are Not independent

Reject H0; Significant; categories are Not independent

Retain H0; Not Significant; categories are independent

Reject H0; Significant; categories are independent

The correct APA format for the result would include

p > .05

p < .05
Answer #1
Applying chi square test of independence:
Expected Ei=row total*column total/grand total A B Total
Good outcome 30.0000 60.0000 90
bad outcome 20.0000 40.0000 60
total 50 100 150
chi square    χ2 =(Oi-Ei)2/Ei A B Total
Good outcome 0.833 0.417 1.2500
bad outcome 1.250 0.625 1.8750
total 2.0833 1.0417 3.1250
test statistic X2= 3.1250

The expected frequency, fe , for the Team A X Good Outcome category = 30

The Chi Square test statistic = 3.125

degree of freedom(df) =(rows-1)*(columns-1)=1

Chi Square critical boundary = 3.841

Retain H0; Not Significant; categories are independent

The correct APA format for the result would include p >0.05